Terry Watkins can row about 10.6 kilometers in hour downstream and 6.8 kilometers upstream in 1 hour. Find how fast he can row in still water, and find the speed of the current.
Answer by mananth(16949)

let speed of boat in still water be x mph
speed of current be y mph
time = d/r
Downstream time
10.6/(x+y)=1
x+y =10.6........................1
...
Upstream time
6.8/(x-y)=1
x-y =6.8........................2
add the two equations (1)&(2)
2x=17.4
/2
x=17.4/2
x= 8.7 mph speed of boat in still water
y=1.9 mph speed of current
...
m.ananth@hotmail.ca